# Partial information decomposition as information bottleneck

### v1

@@ III Redundancy bottleneck
In this section, we introduce the redundancy bottleneck (RB) and illustrate it with examples.
@@ III.1 Reformulation of Blackwell redundancy
We first reformulate Blackwell redundancy (7) in terms of a different optimization problem.
We begin with a few preliminaries. We assume that we are provided with the marginal distribution $p_{Y}$ of the target random variable $Y$, as well as $n$ source channels $p_{X_{1}|Y},\dots,p_{X_{n}|Y}$. Without loss of generality, we assume that $p_{Y}$ has full support. We use calligraphic letters (like $\mathcal{Y}$ and $\mathcal{X}_{s}$) to indicate the set of outcomes of random variables (like $Y$ and $X_{s}$). For simplicity, we generally use notation appropriate for discrete-valued variables, such as in Eq. (4), though most of our results also apply to continuous-valued variables.
Our reformulation of Blackwell redundancy will make use of the random variable $Y$, along with two additional random variables $S$ and $Z$. The outcomes of $S$ are the indexes of the different sources, $\mathcal{S}=\{1,\dots,n\}$. The set of outcomes of $Z$ is the union of the outcomes of the individual sources, $\mathcal{Z}=\bigcup_{s=1}^{n}\mathcal{X}_{s}$. For example, if there are two sources, one binary-valued ($\mathcal{X}_{1}=\{0,1\}$) and ternary-valued ($\mathcal{X}_{2}=\{0,1,2\}$), then the outcomes of $Z$ would be $\mathcal{Z}=\{0,1\}\cup\{0,1,2\}=\{0,1,2\}$. The joint probability distribution over $(Y,S,Z)$ is defined as
@@
In words, $y$ is drawn from the marginal $p_{Y}$, the source $s$ is then drawn independently from the distribution $\nu_{S}$, and finally $z$ is drawn from the channel $p_{X_{s}|Y}(z|y)$ corresponding to source $s$. In this way, the channels corresponding to the $n$ sources ($p_{X_{1}|Y},\dots,p_{X_{n}|Y}$) are combined into a single conditional distribution $p_{Z|SY}$.
We treat the distribution $\nu_{S}$ as an arbitrary fixed parameter. As we will see below, different choices of $\nu_{S}$ will cause the different sources to be weighed differently in the computation of the RB. Except where otherwise noted, we make no assumptions about the distribution $\nu_{S}$ except that it has full support.
Note that $Y$ and $S$ are independent under the distribution defined in Eq. (8), so
@@
In addition to $Y$, $Z$ and $S$, we introduce another random variable $Q$. This random variable obeys the Markov condition $Q-(Z,S)-Y$, which ensures that $Q$ does not contain any information about $Y$ that is not contained in the joint outcome of $Z$ and $S$. The full joint distribution over $(Y,S,Z,Q)$ is
$p_{YSZQ}(s,x,y)=p_{YSZ}(y,s,z)p_{Q|SZ}(q|s,x).$ (10)
@@
Our first result shows that Blackwell redundancy can be equivalently expressed via a constrained optimization problem. Here the optimization is over bottleneck random variables $Q$, i.e., over conditional distributions $p_{Q|SZ}$ in Eq. (10).
@@
$I_{\cap}=\sup_{Q:Q-(Z,S)-Y}I(Q;Y|S)\quad\textrm{where}\quad I(Q;S|Y)=0.$
@@
In Theorem 1, the Blackwell order constraint in Eq. (7) has been replaced by an information-theoretic constraint $I(Q;S|Y)=0$, which states that $Q$ does not provide any information about the identity of source $S$, additional to that already provided by the target $Y$. The objective $I(Q;Y)$ has been replaced by the conditional mutual information $I(Q,Y|S)$. Actually, the objective can be equivalently written in either form, since $I(Q;Y|S)=I(Q;Y)$ given our assumption (see proof of Theorem 1 in the appendix). However, the conditional mutual information form will be useful for further generalization and decomposition, as discussed in the next section.
@@ III.2 Redundancy bottleneck
To relate Blackwell redundancy to IB, we relax the constraint in Theorem 1 by allowing the leakage of $R$ bits of conditional information about the source $S$. This defines the redundancy bottleneck (RB) at compression rate $R$,
We note that for $R>0$, the value of $I_{\text{RB}}(R)$ does depend on the choice of the source distribution $\nu_{S}$.
Eq. (11) is an IB-type problem that involves a tradeoff between prediction $I(Q;Y|S)$ and compression $I(Q;S|Y)$. The prediction term $I(Q;Y|S)$ quantifies the generalized Blackwell redundancy encoded in the bottleneck variable $Q$. The compression term $I(Q;S|Y)$ quantifies the amount of conditional information that the bottleneck variable leaks about the identity of the source. The set of optimal values of $(I(Q;S|Y),I(Q;Y|S))$ define the redundancy bottleneck curve (RB curve) that encodes the overall tradeoff between prediction and compression.
@@
Since $I_{\text{RB}}(R)$ is non-decreasing in $R$, the lowest RB value is achieved in the $R=0$ regime, when it equals the Blackwell redundancy (Theorem 1),
@@
The largest value is achieved as $R\to\infty$, when the compression constraint vanishes. It can be shown that $I(Q;Y|S)\leq I(Z,Y|S)=I(Y;Z,S)$ using the Markov condition $Q-(Z,S)-Y$ and the data processing inequality (see next subsection), and that this upper bound is achieved by using the bottleneck variable $Q=Z$. Combining implies
@@
In addition, for finite-dimensional sources, it suffices to consider finite-dimensional $Q$.
@@
Theorem 3 implies that for finite-dimensional sources, the RB problem (11) involves the maximization of a continuous objective over a compact domain, so the supremum is always achieved by some $Q$  [^1] The cardinality bound for RB is the same as for IB, if we take $X=(Z,S)$ in Eq. (1) [16, 20]. It is larger than the cardinality required for Blackwell redundancy (7), where $|\mathcal{Q}|\leq(\sum_{s}\left|\mathcal{\mathcal{X}}_{s}\right|)-n+1$ suffices [4]..
@@
$F_{\text{RB}}(\beta)=\sup_{Q:Q-(Z,S)-Y}I(Q;Y|S)-\frac{1}{\beta}I(Q;S|Y).$ (14)
The parameter $\beta$ controls the tradeoff between prediction and compression. The $\beta\to 0$ limit corresponds to the $R=0$ regime, in which case Blackwell redundancy is recovered, while the $\beta\to\infty$ limit corresponds to the $R=\infty$ regime, when the compression constraint is removed. The RB Lagrangian (14) is often simpler to optimize than the constrained optimization (11). Moreover, when the RB curve $I_{\text{RB}}(R)$ is strictly concave, there is a one-to-one relationship between the solutions to the two optimization problems (11) and (14). However, when the RB curve is not strictly concave, there is no one-to-one relationship and the usual Lagrangian formulation is insufficient. This can be addressed by optimizing a modified objective that combines prediction and compression in a nonlinear fashion, such as the “exponential Lagrangian” [19],
$F_{\text{RB}}^{\exp}(\beta)=\sup_{Q:Q-(Z,S)-Y}I(Q;Y|S)-\frac{1}{\beta}e^{I(Q;S|Y)}.$ (15)
(See analogous analysis for IB in Ref. [18, 19].)
Figure 1: RB analysis for the UNIQUE gate (Example 1). (a) and (b): Prediction and compression values found by optimizing the RB Lagrangian (14) at different $\beta$ (c) The RB curve shows the tradeoff between optimal compression and prediction values; marker colors correspond to $\beta$ values, as in (a) and (b). All bottleneck variables $Q$ must fall within the accessible grey region. (d) The RB curve is decomposed into contributions from different sources. Here only source 1 provides prediction at any level of compression. Line colors and styles indicate different sources.
@@ III.3 Contributions from different sources
Let $Q$ be an optimal bottleneck variable at rate $R$, so that $I_{\text{RB}}(R)=I(Q;Y|S)$ and $I(Q;S|Y)\leq R$. Then, the RB prediction term can be expressed as the weighted average of prediction contributions from individual sources,
@@
Here we introduced the specific conditional mutual information,
@@
where $D(\cdot\|\cdot)$ is the Kullback-Leibler (KL) divergence. To build intuitions about this decomposition, we use the Markov condition $Q-(Z,S)-Y$ to express the conditional distributions in Eq. (17) as compositions of channels,
@@
Using the data-processing inequality for KL divergence and Eq. (8), we can then write
@@
The last term is simply the mutual information $I(Y;X_{s})$ between the target and source $s$. Thus, the prediction contribution from source $s$ is bounded between 0 and the mutual information provided by that source,
@@
The difference between the mutual information and the actual prediction contribution,
@@
can be understood as the unique information in source $s$. The upper bound in Eq. (18) is achieved in the $R\to\infty$ limit by $Q=Z$, leading to Eq. (13). Conversely, for $R=0$, $p_{Q|Y,S=s}=p_{Q|Y}$ (from $I(Q;S|Y)=0$) and $p_{Q|S=s}=p_{Q}$ (from Eq. (9)), so
@@
Thus, when $R=0$, the prediction contribution from each source is the same, and it is equal to Blackwell redundancy.
@@
Here we introduced the specific conditional mutual information,
@@
Observe that in both decompositions (16) and (19), the source contributions are weighted by the distribution $\nu_{S}(s)$. Thus, the distribution $\nu_{S}$ determines how different sources play into the tradeoff between prediction and compression.
@@
$\delta_{D}(p_{B|Y},p_{C|Y}):=\\ \inf_{\kappa_{B|C}}\sum_{y}p_{Y}(y)D(\kappa_{B|C}\circ p_{C|Y=y}\|p_{B|Y=y}).$ (21)
The weighted deficiency quantifies the degree to which two channels violate the Blackwell order, vanishing when $\kappa_{B|Y}\preceq\kappa_{C|Y}$. To relate the source compression terms (19) to deficiency, observe that $p_{Q|Y,S=s}=p_{Q|Z,S=s}\circ p_{Z|Y,S=s}$ and that $p_{Z|Y,S=s}=p_{X_{s}|Y}$. Given Eq. (20), we then have
$I(Q;S=s|Y)\geq\delta_{D}(p_{X_{s}|Y},p_{Q|Y}).$
Thus, the RB compression constraint in (11) implies an upper bound on the averaged deficiency of $p_{Q|Y}$.
Interestingly, several recent papers have studied the relationship between deficiency and PID redundancy, for the restricted case of two sources [38, 46, 47, 41, 48]. To our knowledge, we provide the first link between deficiency and redundancy for the general case of multiple sources. Note also that previous work considered a slightly different definition of deficiency where the arguments of the KL divergence are reversed. Our definition of deficiency is arguably more natural, since it is more natural to minimize KL over a convex set with respect to the first argument [49].
@@ III.4 Examples
We illustrate our approach using a few examples. For simplicity, in all examples we use a uniform distribution over the sources, $\nu_{S}(s)=1/n$. Numerical results are calculated using the iterative algorithm described in the next section.
Example 1.
We begin by considering a very simple system, called “UNIQUE gate” in the PID literature. Here the target $Y$ is binary and uniformly-distributed, $p_{Y}(y)=1/2$ for $y\in\{0,1\}$. There are two binary-valued sources, $X_{1}$ and $X_{2}$, where the first source is a copy of the target, $p_{X_{1}|Y}(x_{1}|y)=\delta_{x_{1},y}$, while the second source is an independent and uniformly-distributed bit, $p_{X_{2}|Y}(x_{1}|y)=1/2$. Thus, source $X_{1}$ provides 1 bit of information about the target, while $X_{2}$ provides none. The Blackwell redundancy is $I_{\cap}=0$ [4], because it is impossible to extract any information from the sources without revealing that this information came from $X_{1}$.
We perform RB analysis by optimizing the RB Lagrangian $F_{\text{RB}}(\beta)$ (14) at different $\beta$. Figure 1(a) and (b) shows the prediction $I(Q;Y|S)$ and compression $I(Q;S|Y)$ values for the optimal bottleneck variables $Q$. At small $\beta$, prediction converges to Blackwell redundancy, $I(Q;Y|S)=I_{\cap}=0$, and there is complete loss of information about source identity, $I(Q;S|Y)=0$. At larger $\beta$, the prediction approaches the maximum $I(Q;Y|S)=0.5\times I(X_{1};Y)=0.5\text{ bit}$, and compression is $I(Q;S|Y)=I(Z;S|Y)\approx 0.311\text{ bit}$. Figure 1(c) shows the RB curve, showing the overall tradeoff between prediction and compression. Figure 1(d) shows the RB curves for the different sources, showing the source prediction $I(Q;Y|S=s)$ and compression $I(Q;S=s|Y)$ contributions. As expected, only $X_{1}$ contributes to prediction at any level of compression.
@@
Example 2.
We now consider the “AND gate”, another well-known system from the PID literature. There are two independent and uniformly-distributed binary sources $X_{1}$ and $X_{2}$. The target $Y$ is also binary-valued and determined computed via $Y=X_{1}\,\text{AND}\,X_{2}$. Then, $p_{Y}(0)=3/4$ and $p_{Y}(1)=1/4$, and both sources have the same channel,
@@
Because the two sources channels are the same, Blackwell redundancy obeys $I_{\cap}=I(Y;X_{1})=I(Y;X_{2})=0.311$ bits [4]. From Eqs. (12) and (13), we see that $I_{\text{RB}}(R)=I_{\cap}$ across all compression rates. In this system, all information is redundant, so there is no strict tradeoff between prediction and compression. The RB curve (not shown) consists of a single point, $(I(Q;Y|S),I(Q;S|Y))=(0.311,0)$.
Figure 2: RB analysis for the system with 4 binary symmetric channels (Example 3). (a) and (b): Prediction and compression values found by optimizing the RB Lagrangian (14) at different $\beta$. (c) The RB curve shows the tradeoff between optimal compression and prediction values; marker colors correspond to $\beta$ values, as in (a) and (b). All possible bottleneck variables $Q$ must fall within the accessible grey region. (d) The RB curve is decomposed into contributions from different sources. Line colors and styles indicate different sources.
Example 3.
We now consider a more sophisticated example with 4 sources. The target is binary-valued and uniformly-distributed, $p_{Y}(y)=1/2$ for $y\in\{0,1\}$. There are four binary-valued sources, where the conditional distribution of each source $s\in\{1,2,3,4\}$ is a binary symmetric channel with error probability $\epsilon_{s}$:
@@
We take $\epsilon_{1}=\epsilon_{2}=0.1$, $\epsilon_{3}=0.2$, and $\epsilon_{4}=0.5$. Thus, sources $X_{1}$ and $X_{2}$ provide a lot of information about the target, $X_{3}$ provides less information, and while $X_{4}$ is completely independent of the target.
@@
We perform our RB analysis and plot RB prediction values in Figure 2(a) and compression values in Figure 2(b), as found by optimizing the RB Lagrangian at different $\beta$. At small $\beta$, prediction converges to Blackwell redundancy, $I(Q;Y|S)=I_{\cap}=0$, and there is complete loss of information about source identity, $I(Q;S|Y)=0$. At large $\beta$, the prediction is equal to the maximum $I(Z;Y|S)\approx 0.335\text{ bit}$, and compression is equal to $I(Q;S|Y)\approx 0.104\text{ bit}$. The RB curve undergoes some phase transitions as different optimal strategies are uncovered at increasing $\beta$. Figure 1(c) shows the RB curve.

### v2

@@ III Redundancy Bottleneck
In this section, we introduce the redundancy bottleneck (RB) and illustrate it with examples. Generally, we assume that we are provided with the marginal distribution $p_{Y}$ of the target random variable $Y$, as well as $n$ source channels $p_{X_{1}|Y},\dots,p_{X_{n}|Y}$. Without loss of generality, we assume that $p_{Y}$ has full support. We use calligraphic letters (like $\mathcal{Y}$ and $\mathcal{X}_{s}$) to indicate the set of outcomes of random variables (like $Y$ and $X_{s}$). For simplicity, we use notation appropriate for discrete-valued variables, such as in Eq. (4), though most of our results also apply to continuous-valued variables.
@@ III.1 Reformulation of Blackwell Redundancy
We first reformulate Blackwell redundancy (7) in terms of a different optimization problem. Our reformulation will make use of the random variable $Y$, along with two additional random variables, $S$ and $Z$. The outcomes of $S$ are the indexes of the different sources, $\mathcal{S}=\{1,\dots,n\}$. The set of outcomes of $Z$ is the union of the outcomes of the individual sources, $\mathcal{Z}=\bigcup_{s=1}^{n}\mathcal{X}_{s}$. For example, if there are two sources with outcomes $\mathcal{X}_{1}=\{0,1\}$ and $\mathcal{X}_{2}=\{0,1,2\}$, then $\mathcal{S}=\{1,2\}$ and $\mathcal{Z}=\{0,1\}\cup\{0,1,2\}=\{0,1,2\}$. The joint probability distribution over $(Y,S,Z)$ is defined as
@@
In other words, $y$ is drawn from the marginal $p_{Y}$, the source $s$ is then drawn independently from the distribution $\nu_{S}$, and finally $z$ is drawn from the channel $p_{X_{s}|Y}(z|y)$ corresponding to source $s$. In this way, the channels corresponding to the $n$ sources ($p_{X_{1}|Y},\dots,p_{X_{n}|Y}$) are combined into a single conditional distribution $p_{Z|SY}$.
We treat the distribution $\nu_{S}$ as an arbitrary fixed parameter, and except where otherwise noted, we make no assumptions about this distribution except that it has full support. As we will see, different choices of $\nu_{S}$ cause the different sources to be weighed differently in the computation of the RB. We return to the question of how to determine this distribution below.
Note that, under the distribution defined in Eq. (8), $Y$ and $S$ are independent, so
@@
In addition to $Y$, $Z$, and $S$, we introduce another random variable $Q$. This random variable obeys the Markov condition $Q-(Z,S)-Y$, which ensures that $Q$ does not contain any information about $Y$ that is not contained in the joint outcome of $Z$ and $S$. The full joint distribution over $(Y,S,Z,Q)$ is
$p_{YSZQ}(y,s,z,q)=p_{YSZ}(y,s,z)p_{Q|SZ}(q|s,z).$ (10)
@@
The set of joint outcomes of $(S,Z)$ with non-zero probability is the disjoint union of the outcomes of the individual sources. For instance, in the example above with $\mathcal{X}_{1}=\{0,1\}$ and $\mathcal{X}_{2}=\{0,1,2\}$, the set of joint outcomes of $(S,Z)$ with non-zero probability is $\{(1,0),(1,1),(2,0),(2,1),(2,2)\}$. Because $Q$ depends jointly on $S$ and $Z$, our results do not depend on the precise labeling of the source outcomes, e.g., they are the same if $\mathcal{X}_{2}=\{0,1,2\}$ is relabeled as $\mathcal{X}_{2}=\{2,3,4\}$.
Our first result shows that Blackwell redundancy can be equivalently expressed as a constrained optimization problem. Here, the optimization is over bottleneck random variables $Q$, i.e., over conditional distributions $p_{Q|SZ}$ in Eq. (10).
@@
$I_{\cap}=\max_{Q:Q-(Z,S)-Y}I(Q;Y|S)\quad\textrm{where}\quad I(Q;S|Y)=0.$ (11)
@@
In Theorem 1, the Blackwell order constraint in Eq. (7) has been replaced by an information-theoretic constraint $I(Q;S|Y)=0$, which states that $Q$ does not provide any information about the identity of source $S$, additional to that already provided by the target $Y$. The objective $I(Q;Y)$ has been replaced by the conditional mutual information $I(Q;Y|S)$. Actually, the objective can be equivalently written in either form, since $I(Q;Y|S)=I(Q;Y)$ given our assumptions (see the proof of Theorem 1 in the Appendix). However, the conditional mutual information form will be useful for further generalization and decomposition, as discussed in the next sections.
@@ III.2 Redundancy Bottleneck
To relate Blackwell redundancy to the IB, we relax the constraint in Theorem 1 by allowing the leakage of $R$ bits of conditional information about the source $S$. This defines the redundancy bottleneck (RB) at compression rate $R$:
$I_{\text{RB}}(R):=\quad\max_{\mathclap{Q:Q-(Z,S)-Y}}\quad I(Q;Y|S)\quad\textrm{where}\quad I(Q;S|Y)\leq R.$ (12)
We note that, for $R>0$, the value of $I_{\text{RB}}(R)$ does depend on the choice of the source distribution $\nu_{S}$.
Eq. (12) is an IB-type problem that involves a tradeoff between prediction $I(Q;Y|S)$ and compression $I(Q;S|Y)$. The prediction term $I(Q;Y|S)$ quantifies the generalized Blackwell redundancy encoded in the bottleneck variable $Q$. The compression term $I(Q;S|Y)$ quantifies the amount of conditional information that the bottleneck variable leaks about the identity of the source. The set of optimal values of $(I(Q;S|Y),I(Q;Y|S))$ defines the redundancy bottleneck curve (RB curve) that encodes the overall tradeoff between prediction and compression.
@@
Since $I_{\text{RB}}(R)$ is non-decreasing in $R$, the lowest RB value is achieved in the $R=0$ regime, when it equals the Blackwell redundancy (Theorem 1):
@@
The largest value is achieved as $R\to\infty$, when the compression constraint vanishes. It can be shown that $I(Q;Y|S)\leq I(Z;Y|S)=I(Y;Z,S)$ using the Markov condition $Q-(Z,S)-Y$ and the data-processing inequality (see the next subsection). This upper bound is achieved by the bottleneck variable $Q=Z$. Combining implies
@@
Next, we show that, for finite-dimensional sources, it suffices to consider finite-dimensional $Q$. Thus, for finite-dimensional sources, the RB problem (12) involves the maximization of a continuous objective over a compact domain, so the maximum is always achieved by some $Q$. (Conversely, in the more general case of infinite-dimensional sources, it may be necessary to replace $\max$ with $\sup$ in Eq. (12); see Appendix Appendix.)
@@
Interestingly, the cardinality bound for the RB is the same as for the IB if we take $X=(Z,S)$ in Eq. (1) [16, 20]. It is larger than the cardinality required for Blackwell redundancy (7), where $|\mathcal{Q}|\leq(\sum_{s}\left|\mathcal{\mathcal{X}}_{s}\right|)-n+1$ suffices [4].
@@
$F_{\text{RB}}(\beta)=\max_{Q:Q-(Z,S)-Y}I(Q;Y|S)-\frac{1}{\beta}I(Q;S|Y).$ (15)
The parameter $\beta$ controls the tradeoff between prediction and compression. The $\beta\to 0$ limit corresponds to the $R=0$ regime, in which case, Blackwell redundancy is recovered, while the $\beta\to\infty$ limit corresponds to the $R=\infty$ regime, when the compression constraint is removed. The RB Lagrangian (15) is often simpler to optimize than the constrained optimization (12). Moreover, when the RB curve $I_{\text{RB}}(R)$ is strictly concave, there is a one-to-one relationship between the solutions to the two optimization problems (12) and (15). However, when the RB curve is not strictly concave, there is no one-to-one relationship and the usual Lagrangian formulation is insufficient. This can be addressed by optimizing a modified objective that combines prediction and compression in a nonlinear fashion, such as the “exponential Lagrangian” [19]:
$F_{\text{RB}}^{\exp}(\beta)=\max_{Q:Q-(Z,S)-Y}I(Q;Y|S)-\frac{1}{\beta}e^{I(Q;S|Y)}.$ (16)
(See an analogous analysis for IB in Refs. [18, 19].)
@@ III.3 Contributions From Different Sources
Figure 1: RB analysis for the UNIQUE gate (Example 1). (a) Prediction values found by optimizing the RB Lagrangian (15) at different $\beta$. Colored regions indicate contributions from different sources, $\nu_{S}(s)I(Q;Y|S=s)$ from Eq. (17). For this system, only source $X_{1}$ contributes to the prediction. (b) Compression costs found by optimizing the RB Lagrangian at different $\beta$. Colored regions indicate contributions from different sources, $\nu_{S}(s)I(Q;S=s|Y)$ from Eq. (20). (c) The RB curve shows the tradeoff between optimal compression and the prediction values; the marker colors correspond to the $\beta$ values as in (a) and (b). All bottleneck variables $Q$ must fall within the accessible grey region. (d) RB curves for individual sources.
Let $Q$ be an optimal bottleneck variable at rate $R$, so that $I_{\text{RB}}(R)=I(Q;Y|S)$ and $I(Q;S|Y)\leq R$. Then, the RB prediction term can be expressed as the weighted average of the prediction contributions from individual sources:
@@
Here, we introduce the specific conditional mutual information:
@@
where $D(\cdot\|\cdot)$ is the Kullback–Leibler (KL) divergence. To build intuitions about this decomposition, we may use the Markov condition $Q-(Z,S)-Y$ to express the conditional distributions in Eq. (18) as compositions of channels:
@@
Using the data-processing inequality for the KL divergence and Eq. (8), we can then write
@@
The last term is simply the mutual information $I(Y;X_{s})$ between the target and source $s$. Thus, the prediction contribution from source $s$ is bounded between 0 and the mutual information provided by that source:
@@
The difference between the mutual information and the actual prediction contribution:
@@
quantifies the unique information in source $s$. The upper bound in Eq. (19) is achieved in the $R\to\infty$ limit by $Q=Z$, leading to Eq. (14). Conversely, for $R=0$, $p_{Q|Y,S=s}=p_{Q|Y}$ (from $I(Q;S|Y)=0$) and $p_{Q|S=s}=p_{Q}$ (from Eq. (9)), so
@@
Thus, when $R=0$, the prediction contribution from each source is the same, and it is equal to the Blackwell redundancy.
@@
Here, we introduce the specific conditional mutual information:
@@
$\delta_{D}(p_{C|Y},p_{B|Y})$ $:=\min_{\kappa_{B|C}}D(\kappa_{B|C}\circ p_{C|Y}\|p_{B|Y}).$ (22)
This measure quantifies the degree to which two channels violate the Blackwell order, vanishing when $\kappa_{B|Y}\preceq\kappa_{C|Y}$. To relate the source compression terms (20) to deficiency, observe that $p_{Q|Y,S=s}=p_{Q|Z,S=s}\circ p_{Z|Y,S=s}$ and that $p_{Z|Y,S=s}=p_{X_{s}|Y}$. Given Eq. (21), we then have
$I(Q;S=s|Y)\geq\delta_{D}(p_{X_{s}|Y},p_{Q|Y}).$ (23)
Thus, each source compression term is lower bounded by the deficiency between the source channel $p_{X_{s}|Y}$ and the bottleneck channel $p_{Q|Y}$. Furthermore, the compression constraint in the RB optimization problem (12) sets an upper bound on the deficiency of $p_{Q|Y}$ averaged across all sources.
Interestingly, several recent papers have studied the relationship between deficiency and PID redundancy in the restricted case of two sources [38, 45, 46, 41, 47]. To our knowledge, we provide the first link between deficiency and redundancy for the general case of multiple sources. Note also that previous work considered a slightly different definition of deficiency where the arguments of the KL divergence are reversed. Our definition of deficiency is arguably more natural, since it is more natural to minimize the KL divergence over a convex set with respect to the first argument [48].
Finally, observe that, in both decompositions (17) and (20), the source contributions are weighted by the distribution $\nu_{S}(s)$. Thus, the distribution $\nu_{S}$ determines how different sources play into the tradeoff between prediction and compression. In many cases, $\nu_{S}$ can be chosen as the uniform distribution. However, other choices of $\nu_{S}$ may be more natural in other situations. For example, in a neuroscience context where different sources correspond to different brain regions, $\nu_{S}(s)$ could represent the proportion of metabolic cost or neural volume assigned to region $s$. Alternatively, when different sources represent mutually exclusive conditions, as in the age group example mentioned at the end of Section II, $\nu_{S}(s)$ might represent the frequency of condition $s$ found in the data. Finally, it may be possible to set $\nu_{S}$ in an “adversarial” manner so as to maximize the resulting value of $I_{\text{RB}}(R)$ in Eq. (12). We leave the exploration of this adversarial approach for future work.
@@ III.4 Examples
We illustrate our approach using a few examples. For simplicity, in all examples, we use a uniform distribution over the sources, $\nu_{S}(s)=1/n$. The numerical results are calculated using the iterative algorithm described in the next section.
Example 1. We begin by considering a very simple system, called the “UNIQUE gate” in the PID literature. Here, the target $Y$ is binary and uniformly-distributed, $p_{Y}(y)=1/2$ for $y\in\{0,1\}$. There are two binary-valued sources, $X_{1}$ and $X_{2}$, where the first source is a copy of the target, $p_{X_{1}|Y}(x_{1}|y)=\delta_{x_{1},y}$, while the second source is an independent and uniformly-distributed bit, $p_{X_{2}|Y}(x_{1}|y)=1/2$. Thus, source $X_{1}$ provides 1 bit of information about the target, while $X_{2}$ provides none. The Blackwell redundancy is $I_{\cap}=0$ [4], because it is impossible to extract any information from the sources without revealing that this information came from $X_{1}$.
Figure 2: RB analysis for the system with 4 binary symmetric channels (Example 3). (a) and (b) Prediction and compression values found by optimizing the RB Lagrangian (15) at different $\beta$. Contributions from individual sources are shown as shaded regions. (c) The RB curve shows the tradeoff between optimal compression and prediction values; marker colors correspond to the $\beta$ values as in (a) and (b). (d) RB curves for individual sources.
We performed RB analysis by optimizing the RB Lagrangian $F_{\text{RB}}(\beta)$ (15) at different $\beta$. Figure 1(a) and (b) show the prediction $I(Q;Y|S)$ and compression $I(Q;S|Y)$ values for the optimal bottleneck variables $Q$. At small $\beta$, the prediction converges to the Blackwell redundancy, $I(Q;Y|S)=I_{\cap}=0$, and there is complete loss of information about source identity, $I(Q;S|Y)=0$. At larger $\beta$, the prediction approaches the maximum $I(Q;Y|S)=0.5\times I(X_{1};Y)=0.5\text{ bit}$, and compression approaches $I(Q;S|Y)=I(Z;S|Y)\approx 0.311\text{ bit}$. Figure 1(c) shows the RB curve, illustrating the overall tradeoff between prediction and compression.
In the shaded regions of Figure 1(a) and (b), we show the additive contributions to the prediction and compression terms from the individual sources, $\nu_{S}(s)I(Q;Y|S=s)$ from Eq. (17) and $\nu_{S}(s)I(Q;S=s|Y)$ from Eq. (20), respectively. We also show the resulting RB curves for individual sources in Figure 1(d). As expected, only source $X_{1}$ contributes to the prediction at any level of compression.
@@
Example 2. We now consider the “AND gate”, another well-known system from the PID literature. There are two independent and uniformly distributed binary sources, $X_{1}$ and $X_{2}$. The target $Y$ is also binary-valued and determined via $Y=X_{1}\,\text{AND}\,X_{2}$. Then, $p_{Y}(0)=3/4$ and $p_{Y}(1)=1/4$, and both sources have the same channel:
@@
Because the two source channels are the same, the Blackwell redundancy obeys $I_{\cap}=I(Y;X_{1})=I(Y;X_{2})=0.311$ bits [4]. From Eqs. (13) and (14), we see that $I_{\text{RB}}(R)=I_{\cap}$ across all compression rates. In this system, all information provided by the sources is redundant, so there is no strict tradeoff between prediction and compression. The RB curve (not shown) consists of a single point, $(I(Q;Y|S),I(Q;S|Y))=(0.311,0)$.
Figure 3: RB analysis for the system with a 3-spin target (Example 4). (a) and (b) Prediction and compression values found by optimizing the RB Lagrangian (15) at different $\beta$. Contributions from individual sources are shown as shaded regions. (c) The RB curve shows the tradeoff between optimal compression and prediction values; marker colors correspond to the $\beta$ values as in (a) and (b). (d) RB curves for individual sources.
Example 3. We now consider a more sophisticated example with four sources. The target is binary-valued and uniformly distributed, $p_{Y}(y)=1/2$ for $y\in\{0,1\}$. There are four binary-valued sources, where the conditional distribution of each source $s\in\{1,2,3,4\}$ is a binary symmetric channel with error probability $\epsilon_{s}$:
@@
We take $\epsilon_{1}=\epsilon_{2}=0.1$, $\epsilon_{3}=0.2$, and $\epsilon_{4}=0.5$. Thus, sources $X_{1}$ and $X_{2}$ provide most information about the target; $X_{3}$ provides less information; $X_{4}$ is completely independent of the target.
@@
We performed our RB analysis and plot the RB prediction values in Figure 2(a) and the compression values in Figure 2(b), as found by optimizing the RB Lagrangian at different $\beta$. At small $\beta$, the prediction converges to the Blackwell redundancy, $I(Q;Y|S)=I_{\cap}=0$, and there is complete loss of information about source identity, $I(Q;S|Y)=0$. At large $\beta$, the prediction is equal to the maximum $I(Z;Y|S)\approx 0.335\text{ bit}$, and compression is equal to $I(Q;S|Y)\approx 0.104\text{ bit}$. Figure 2(c) shows the RB curve.
In Figure 2(a) and (b), we show the additive contributions to the prediction and compression terms from the individual sources, $\nu_{S}(s)I(Q;Y|S=s)$ and $\nu_{S}(s)I(Q;S=s|Y)$, respectively, as shaded regions. We also show the resulting RB curves for individual sources in Figure 2(d).